OpenStudy (anonymous):

let f(x)=2x^4 - 4x^2 find all c in the interval (-3, 3) such that f'(c) = 0. PLEASE HELP AND SHOW ALL STEPS!

OpenStudy (anonymous):

f'(x) = 2x^2 - 8x = 0 8x = 2x^2 4 = 2x 2 = x

OpenStudy (jamesj):

There are a couple of errors in what Diarmuid has written, beginning with the derivative. Can you calculate the derivative yourself?

OpenStudy (anonymous):

Damn. sorry. I should have read the question more carefully f'(x) = 8x^3 -8x I was wondering about the wording when it looked like there was only one minimum. 0 = 8x^3 -8x 1 = x^2 x = + - 1 and of course 0 (should there be a mathematical way of getting the zero?) so points are (-1,0), (0,0), 1,0)